# SFormer-OR: Single-Stage Bimodal Transformer for Scene Graph Generation in OR

### v1

@@ 5 Discussion
Scene graph generation (SGG) in the operating room (OR) is essential for monitoring the surgical process, optimizing team collaboration, and improving workflow efficiency. Compared to reviewing surveillance records, generating scene graphs can highlight specific events of the surgical process and save the time cost of identifying abnormal events. However, it is challenging to generate scene graphs directly from multimodal data streams. Existing OR-SGG approaches [6, 5] first detect human poses and objects, which are then used as inputs to predict scene graphs via GCNs. Although this multi-stage architecture reaches good performance, it is complicated and difficult to deploy in real clinical practices. Driven by the above motivations, this paper aims to infer SGG in ORs using 2D multi-view images and 3D point cloud data without intermediate steps of processing.
@@
We propose a novel single-stage multi-view bimodal transformer (SFormer-OR) model for SGG in ORs. In practice, our model is constructed primarily based on a 2D SGG baseline augmented by point cloud samples. In complex surgical procedures, leveraging point clouds to enhance geometric perception is crucial for prediction robustness. Besides, the data streams from multiple cameras are also essential for monitoring specific actions of the head surgeon and overall scene recognition. Therefore, our single-stage model includes a View-Sync Transfusion (VST) module for synthesizing multi-view semantic features and a Geometry-Visual Cohesion (GVC) operation for injecting structural cues extracted from point cloud features into synergic features. To generate scene graphs in an end-to-end manner, we embed a relation-sensitive transformer in SFormer-OR, which utilizes unified features and entity-aware embeddings to predict subject-object relations.
@@
As the only benchmark for OR SGG, the 4D-OR dataset [6] contains a variety of interaction relations (e.g., Assisting, Drilling, and Touching) and spatial structural relations (e.g., Close-to and Lying-on). Unlike general 3D SGG datasets [7] that focus primarily on spatial relations, 4D-OR requires more precise prediction of human-human or human-object interactions [38, 39, 40]. In this regard, our method is dedicated to the fine-grained extraction and fusion of multi-view semantic features, especially with our VST scheme and relational trait priors. As shown in Table 1, SFormer-OR achieves superior performance in predicting relationships initiated by surgeons, especially some subtle actions, such as Cementing, Cleaning, Cutting, and Sawing. Such capability to robustly predict subtle relationships benefits our model to be applied not only to knee surgery but also to be generalized to other types of surgery. We believe our method can inspire more investigations on single-stage scene graph generation in operating rooms, further boosting the efficiency of surgical process analysis and promoting the development of surgical intelligence.
@@ 6 Conclusion
We present a novel single-stage multi-view bimodal transformer (SFormer-OR) method for scene graph generation in operating rooms. Specifically, we design a View-Sync Transfusion (VST) scheme to synthesize scene information from multiple viewpoints and a Geometry-Visual Cohesion (GVC) operation to implicitly merge 2D semantic features with 3D point cloud geometric cues. Furthermore, a relation-sensitive transformer is proposed to directly predict scene graphs via dynamic relation queries and blend multimodal features. The promising performance on the 4D-OR dataset demonstrates the effectiveness of our approach. More importantly, compared to existing multi-stage OR-SGG models, our single-stage end-to-end framework can generate scene graphs directly from surgical process content, greatly facilitating real-life applications in clinical surgical scenes.

### v2

@@ 5 Discussion
Scene graph generation (SGG) in the operating room (OR) is essential for monitoring the surgical process, optimizing team collaboration, and improving workflow efficiency. Compared to reviewing surveillance records, generating scene graphs can highlight specific events of the surgical process and save the time cost of identifying abnormal events. However, it is challenging to generate scene graphs directly from multimodal data streams. Existing OR-SGG approaches [6, 5] first detect human poses and objects, which are then used as inputs to predict scene graphs via GCNs. Although this multi-stage architecture yields good performance, it requires stage-by-stage fine-tuning, which affects the overall effectiveness of end-to-end training of the model. Additionally, the multi-stage model is more cumbersome to deploy in real clinical practices, making it difficult to land in real-world surgery. Driven by the above motivations, this paper aims to infer scene graphs in ORs using 2D multi-view images and 3D point cloud data without intermediate steps of processing.
@@
We propose a novel single-stage multi-view bi-modal transformer (S2Former-OR) model for SGG in ORs. In practice, our model is constructed primarily based on a 2D SGG baseline augmented by point cloud samples. In complex surgical procedures, leveraging point clouds to enhance geometric perception is crucial for prediction robustness. Besides, the data streams from multiple cameras are also essential for monitoring specific actions of the head surgeon and overall scene recognition. Therefore, our single-stage model includes a View-Sync Transfusion (VST) module for synthesizing multi-view semantic features and a Geometry-Visual Cohesion (GVC) operation for injecting structural cues extracted from point cloud features into synergic features. To generate scene graphs in an end-to-end manner, we embed a relation-sensitive transformer in S2Former-OR, which utilizes unified features and entity-aware embeddings to predict subject-object relations.
@@
As the only benchmark for OR SGG, the 4D-OR dataset [6] contains a variety of interaction relations (e.g., Assisting, Drilling, and Touching) and spatial structural relations (e.g., Close-to and Lying-on). Unlike general 3D SGG datasets [7] that focus primarily on spatial relations, 4D-OR requires more precise prediction of human-human or human-object interactions [41, 42, 43]. In this regard, our method is dedicated to the fine-grained extraction and fusion of multi-view semantic features, especially with our VST scheme and relational trait priors. As shown in Table 1, S2Former-OR achieves superior performance in predicting relationships initiated by surgeons, especially some subtle actions, such as Cementing, Cleaning, Cutting, and Sawing. Such capability to robustly predict subtle relationships benefits our model to be applied not only to knee surgery but also to be generalized to other types of surgery. We believe our method can inspire more investigations on single-stage scene graph generation in operating rooms, further boosting the efficiency of surgical process analysis and promoting the development of surgical intelligence.
@@ 6 Conclusion
We present a novel single-stage multi-view bi-modal transformer (S2Former-OR) method for scene graph generation in operating rooms. Specifically, we design a View-Sync Transfusion (VST) scheme to synthesize scene information from multiple viewpoints and a Geometry-Visual Cohesion (GVC) operation to implicitly merge 2D semantic features with 3D point cloud geometric cues. Furthermore, a relation-sensitive transformer is proposed to directly predict scene graphs via dynamic relation queries and blend multimodal features. The promising performance on the 4D-OR dataset demonstrates the effectiveness of our approach. More importantly, compared to existing multi-stage OR-SGG models, our single-stage end-to-end framework can generate scene graphs directly from surgical process content, greatly facilitating real-life applications in clinical surgical scenes.